A researcher used a developed problem solving test to randomly select 50 Grade 6 pupils. In


this sample, and . The mean and the standard deviation of the population used in


the standardization of the test were 75 and 15, respectively. Use the 95% confidence level

1
Expert's answer
2022-06-06T06:33:39-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=75"

"H_1:\mu\not=75"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z:|z|>1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{80-75}{15/\sqrt{6}}=0.8165"

Since it is observed that "z=0.8165<1.96=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=2P(z>0.8165)= 0.414214," and since "p=0.414214>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than 75, at the "\alpha = 0.05" significance level.